# Improving the statistical efficiency of cross-conformal prediction

### v1

@@ Abstract
Vovk, (2015) introduced cross-conformal prediction, a modification of split conformal designed to improve the width of prediction sets. The method, when trained with a miscoverage rate equal to $\alpha$ and $n\gg K$, ensures a marginal coverage of at least $1-2\alpha-2(1-\alpha)(K-1)/(n+K)$, where $n$ is the number of observations and $K$ denotes the number of folds. A simple modification of the method achieves coverage of at least $1-2\alpha$. In this work, we propose new variants of both methods that yield smaller prediction sets without compromising the latter theoretical guarantee. The proposed methods are based on recent results deriving more statistically efficient combination of p-values that leverage exchangeability and randomization. Simulations confirm the theoretical findings and bring out some important tradeoffs.
@@ 1 Introduction
Conformal prediction has emerged as a general and versatile framework for constructing prediction sets in regression and classification tasks (Shafer and Vovk,, 2008). Unlike traditional methods, which often depend on rigid distributional assumptions, conformal prediction transforms point predictions from any prediction (or black-box) algorithm into prediction sets that guarantee valid finite-sample marginal coverage. Originally introduced by Vovk et al., (2005), it has become increasingly influential, with numerous methods and extensions being proposed since its introduction.
@@
In Section 2 we illustrate the problem setup and related work. In Section 3 cross-conformal prediction is described while the new methods and results are presented in Section 4. Section 5 presents some empirical results. In particular, Section 5.1 contains some simulation results, while an application to a real-world dataset is presented in Section 5.2.
@@ 2 Problem setup and related work
Assume we have independent and identically distributed (iid) training samples $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y},\,i=1,\dots,n,$ drawn from a probability distribution $Q$, where $\mathcal{X}$ represents the feature space and $\mathcal{Y}$ the response space. Using these training data, our goal is to obtain a prediction set for the response variable $Y_{n+1}$ based on the covariates $X_{n+1}$, under the assumption that the test pair $(X_{n+1},Y_{n+1})$ is independently sampled from the same distribution $Q$. A typical scenario involves applying a prediction algorithm to the training data in order to find a prediction for the response value. In particular, let $\hat{\mu}:\mathcal{X}\to\mathcal{Y}^{\prime}$ be a regression function obtained applying an algorithm $\mathcal{A}$ to the training points, where $\mathcal{Y^{\prime}}$ is the prediction space (in regression problems we usually have $\mathcal{X}=\mathbb{R}^{p}$ and $\mathcal{Y}=\mathcal{Y^{\prime}}=\mathbb{R})$. Formally, $\mathcal{A}$ is a mapping from $\cup_{d\geq 1}(\mathcal{X}\times\mathcal{Y})^{d}$ (the set of all possible training datasets of any size $d\geq 1$), to the space of functions $\mathcal{X}\to\mathcal{Y}^{\prime}$. Starting from the regression function $\hat{\mu}$, we aim to construct a prediction set $\hat{C}(X_{n+1})$ that contains the point $Y_{n+1}$ with high probability. Since no assumptions are made about the distribution $Q$, the method is said to be distribution-free.
@@
where $\pi$ is any permutation of the indices $[d]:=\{1,\dots,d\}$ and $\mathcal{D}^{\pi}$ refer to the dataset whose elements are permuted by $\pi$. For instance, when considering residual scores $|y-\hat{\mu}(x)|$ in a regression problem, the symmetry of the score function implies that the prediction algorithm is symmetric, meaning that $\mathcal{A}(\mathcal{D})=\mathcal{A}(\mathcal{D}^{\pi})$. In addition, we denote the dataset containing the observations in the set $I$ as $\mathcal{D}_{I}=(Z_{i}:i\in I)$.
@@ 2.1 Related work
As outlined in Section 1, full conformal prediction was introduced in Vovk et al., (2005) and other influential works in the conformal prediction framework include Lei et al., (2018), Romano et al., (2019) and Barber et al., (2021). Extensions of these methods are proposed in Kim et al., (2020) and Gupta et al., (2022). Our work is based on the cross-conformal prediction method introduced in Vovk, (2015) and later extended in Vovk et al., (2018). We refer to Fontana et al., (2023) and Angelopoulos and Bates, (2023) for an overview of conformal prediction and its extensions.
@@
In the framework of conformal prediction, the combination (or ensembling) of p-values is used in Carlsson et al., (2014), Toccaceli and Gammerman, (2017) and Linusson et al., (2017). Their empirical results indicate that Fisher’s method is not a valid rule for combining p-values obtained from different splits or algorithms, whereas using rules valid under arbitrary dependence tends to be generally conservative. In particular, Linusson et al., (2017) provides some intuitions suggesting that the empirical coverage of cross-conformal prediction depends on the degree of dependence between the conformal p-values (that it is strictly related to the stability of the underlying prediction algorithm). In a similar spirit, the solutions in Cherubin, (2019) and Solari and Djordjilović, (2022) aim to combine dependent conformal prediction sets (rather than p-values) derived from different random splits or prediction algorithms. Gasparin and Ramdas, (2024) extended this approach by incorporating a weighting system and including randomization.
@@ 3.1 Cross-conformal prediction
Cross-conformal prediction, introduced by Vovk, (2015), is a method to obtain distribution-free prediction sets. It can be considered as a combination of split conformal prediction (see Appendix A) and cross-validation. It works as follows: data are divided into $K$ disjoint subsets (or folds) $I_{1},\dots,I_{K}$ of size $m=n/K$. The (cross-validation) scores are defined as:
@@
This implies that the statement in (4) can be proved by combining the results in (6) and (7).
@@ 3.2 Modified cross-conformal prediction
Since p-values take discrete values, in order to avoid having noninformative sets identical to $\mathcal{Y}$, the inequality $1<\alpha(m+1)$ must hold. A slight improvement can be obtained using randomized p-values $P_{1}(Y_{n+1};\tau),\dots,P_{K}(Y_{n+1};\tau)$ defined by
$\begin{split}&P_{k}(y;\tau)=\frac{\tau+\sum_{i\in I_{k}}\tau\mathds{1}\left\{s\left((X_{n+1},y);\mathcal{D}_{[n]\setminus I_{k}}\right)=S_{i}^{\mathrm{CV}}\right\}+\mathds{1}\left\{s\left((X_{n+1},y);\mathcal{D}_{[n]\setminus I_{k}}\right)<S_{i}^{\mathrm{CV}}\right\}}{m+1},\end{split}$ (9)
@@ 4 New variants of cross-conformal prediction
In this section, we improve the prediction set in (8) using recent results regarding the combination of p-values. In particular, the combination rules that will be used are more powerful than the combinations valid under arbitrary dependence of the p-values. The results are obtained in a completely general manner and do not require the use of expensive computational procedures (Carlsson et al.,, 2014) or the use of specific models (Boström et al.,, 2017).
@@ 4.1 Exchangeable modified cross-conformal prediction
Let $P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1})$ be the (cross-conformal) p-values obtained using data $Z_{i}=(X_{i},Y_{i}),\,i=[n+1],$ then $P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1})$ are exchangeable, meaning that
$\mathbf{P}\stackrel{{\scriptstyle d}}{{=}}\mathbf{P}^{\pi},$
where =$\stackrel{{\scriptstyle d}}{{=}}$ represents equality in distribution, $\mathbf{P}=(P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1}))$, $\mathbf{P}^{\pi}=(P_{\pi(1)}(Y_{n+1}),\dots,P_{\pi(K)}(Y_{n+1}))$ and $\pi:[K]\to[K]$ is any permutation of the indices.
@@
It holds that $\hat{C}_{n,K,\alpha}^{\mathrm{e\text{-}mod\text{-}cross}}(X_{n+1})\subseteq\hat{C}_{n,K,\alpha}^{\mathrm{mod\text{-}cross}}(X_{n+1})$. In addition,
@@
Once $K$ exchangeable (or more generally dependent) p-values are obtained, there are several methods to combine them. The proposed solution is to use the minimum (over $\ell$) of the mean obtained using the first $\ell$ p-values, which is related to the valid combination rule “twice the average” used by Vovk et al., (2018); Vovk et al., 2022a to prove the coverage guarantee of cross-conformal prediction. However, similar results apply to other merging functions like quantiles (for example “twice the median” is also a valid combination rule) and generalized averages (e.g., geometric mean or harmonic mean).
@@ 4.2 Randomized modified cross-conformal prediction
In the previous paragraph, we leveraged the exchangeability of p-values to obtain a smaller set. In this section, we move in a different direction and improve the set (8) using a simple “randomization trick” (introducing a uniform random variable). As before, the improvement does not alter the marginal validity of the set, but the new result is obtained in a different way. Although randomization is avoided in some statistical applications due to the extra randomness it introduces, in this case, it does not pose a major issue. Indeed, cross-conformal prediction is, by definition, a randomized method. More precisely, data are randomly divided into $K$ different subsets in the first step, which means that the procedure inherently includes randomness (see Remark 4.8 for further discussion).
@@
$\begin{split}\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha}&(X_{n+1})=\Bigg{\{}y\in\mathcal{Y}:\frac{1}{2-U}\frac{1}{K}\sum_{k=1}^{K}P_{k}(y)>\alpha\Bigg{\}},\end{split}$ (12)
@@
It holds that $\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$. In addition,
@@ 4.3 Exchangeable and randomized modified cross-conformal prediction
$\begin{split}&\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})=\Bigg{\{}y\in\mathcal{Y}:\min\left\{\frac{1}{2-U}P_{1}(y),\min_{\ell\in[K]}\frac{1}{\ell}\sum_{k=1}^{\ell}P_{k}(y)\right\}>\alpha\Bigg{\}},\end{split}$ (14)
@@
It holds that $\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{e\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$. In addition,
@@
The set in (14) can be considered an improvement of the set described in (10) but not of the (randomized) set in (12), since only the first p-value of the sequence is randomized (see Table 4 in Appendix E for an example).
@@
A direct use of randomization is present in both procedures described in Section 4.2 and Section 4.3. The use of randomization is often avoided in statistical methods, as it can pose challenges to the reproducibility of results. Clearly, randomization becomes problematic when a human is in the loop and runs the procedure multiple times until the desired result is achieved (for example, in the described cases, one can sample $U$ many times until it reaches a value close to zero). Some recommendations aimed at solving this problem are proposed, for example, in Ramdas and Manole, (2025, Section 10). Actually, in the data pipeline of split and cross-conformal prediction methods, randomization comes into play in different parts: by default in the division of data into their respective folds; to smoothen p-values as described in (9); and potentially to improve the conditional coverage as described in Hore and Barber, (2024). In particular, there exists a trade-off between reproducibility and statistical efficiency, and it is not always evident which should be prioritized. In other words, randomized procedures tend to be more efficient than standard procedures but may lack in terms of reproducibility, and vice versa. For instance, our methods may be particularly well-suited in industrial settings, where hundreds or thousands of predictions are made daily, and efficiency may be more important.

### v2

@@ Abstract
Vovk, (2015) introduced cross-conformal prediction, a modification of split conformal designed to improve the width of prediction sets. The method, when trained with a miscoverage rate equal to $\alpha$ and $n\gg K$, ensures a marginal coverage of at least $1-2\alpha-2(1-\alpha)(K-1)/(n+K)$, where $n$ is the number of observations and $K$ denotes the number of folds. A simple modification of the method achieves coverage of at least $1-2\alpha$. In this work, we propose new variants of both methods that yield smaller prediction sets without compromising the latter theoretical guarantees. The proposed methods are based on recent results deriving more statistically efficient combination of p-values that leverage exchangeability and randomization. Simulations confirm the theoretical findings and bring out some important tradeoffs.
@@ 1 Introduction
Conformal prediction has emerged as a general and versatile framework for constructing prediction sets in regression and classification tasks (Shafer and Vovk,, 2008). Unlike traditional methods, which often depend on rigid distributional assumptions, conformal prediction transforms point predictions from any prediction (or black-box) algorithm into prediction sets that guarantee valid finite-sample marginal coverage. Originally introduced by Saunders et al., (1999), it has become increasingly influential, with numerous methods and extensions being proposed since its introduction.
@@
The structure of the paper is as follows. In Section 2 we illustrate the problem setup and related work. In Section 3 cross-conformal prediction is described while the new methods and results are presented in Section 4. Section 5 presents some empirical results. In particular, Section 5.1 contains some simulation results, while an application to a real-world dataset is presented in Section 5.2.
@@ 2 Problem setup and related work
Assume we have independent and identically distributed (iid) training samples $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathcal{Y},\,i=1,\dots,n,$ drawn from a probability distribution $Q$, where $\mathcal{X}$ represents the feature space and $\mathcal{Y}$ the response space. Using these training data, our goal is to obtain a prediction set for the response variable $Y_{n+1}$ based on the covariates $X_{n+1}$, under the assumption that the test pair $(X_{n+1},Y_{n+1})$ is independently sampled from the same distribution $Q$. In what follows, the results will be shown to hold more generally under the assumptions of exchangeability of the $n+1$ data points with the iid assumption as a special case. A typical scenario involves applying a prediction algorithm to the training data in order to find a prediction for the response value. In particular, let $\hat{\mu}:\mathcal{X}\to\mathcal{Y}^{\prime}$ be a regression function obtained by applying an algorithm $\mathcal{A}$ to the training points, where $\mathcal{Y^{\prime}}$ is the prediction space (in regression problems we usually have $\mathcal{X}=\mathbb{R}^{p}$ and $\mathcal{Y}=\mathcal{Y^{\prime}}=\mathbb{R})$. Formally, $\mathcal{A}$ is a mapping from $\cup_{d\geq 1}(\mathcal{X}\times\mathcal{Y})^{d}$ (the set of all possible training datasets of any size $d\geq 1$), to the space of functions $\mathcal{X}\to\mathcal{Y}^{\prime}$. Starting from the regression function $\hat{\mu}$, we aim to construct a prediction set $\hat{C}(X_{n+1})$ that contains the point $Y_{n+1}$ with high probability. Since no assumptions are made about the distribution $Q$, the method is said to be distribution-free.
@@
where $\pi$ is any permutation of the indices $[d]:=\{1,\dots,d\}$ and $\mathcal{D}^{\pi}$ refer to the dataset whose elements are permuted by $\pi$. For example, when considering residual scores $|y-\hat{\mu}(x)|$ in a regression problem, the symmetry of the score function is satisfied if the prediction algorithm is symmetric, which means that $\mathcal{A}(\mathcal{D})=\mathcal{A}(\mathcal{D}^{\pi})$. In addition, we denote the dataset containing the observations in the set $I$ as $\mathcal{D}_{I}=(Z_{i}:i\in I)$.
@@ 2.1 Related work
As outlined in the Introduction, conformal prediction was first introduced and formalized by Saunders et al., (1999) and Vovk et al., (2005). Several influential contributions to the framework include the works of Lei et al., (2018), Romano et al., (2019), and Barber et al., (2021). Extensions and generalizations of the methods have been proposed by Kim et al., (2020) and Gupta et al., (2022), among others. Other works extend conformal prediction to settings where the standard assumptions may not hold, such as Tibshirani et al., (2019), Prinster et al., (2022), Barber et al., (2023) and Stutz et al., (2023). Our work is based on the cross-conformal prediction method introduced in Vovk, (2015) and later extended in Vovk et al., (2018). We refer to Fontana et al., (2023) and Angelopoulos and Bates, (2023) for an overview of conformal prediction and its extensions.
@@
In the framework of conformal prediction, the combination (or ensembling) of p-values is used in Carlsson et al., (2014), Toccaceli and Gammerman, (2017) and Linusson et al., (2017). Their empirical results indicate that Fisher’s method is not a valid rule for combining p-values obtained from different splits or algorithms, whereas using rules valid under arbitrary dependence tends to be generally conservative. In particular, Linusson et al., (2017) provides some intuitions suggesting that the empirical coverage of cross-conformal prediction depends on the degree of dependence between the conformal p-values (that it is strictly related to the stability of the underlying prediction algorithm). In a similar spirit, the solutions in Cherubin, (2019) and Solari and Djordjilović, (2022) aim to combine dependent conformal prediction sets (rather than p-values) derived from different random splits or prediction algorithms. In particular, their approach relies on a majority vote strategy. Gasparin and Ramdas, (2024) further extended their approach by introducing a weighting system and incorporating randomization. Other works aim to select or combine conformal prediction sets; see for example Yang and Kuchibhotla, (2024) and Liang et al., (2024).
@@ 3.1 Cross-conformal prediction
Cross-conformal prediction, introduced by Vovk, (2015), is a method to obtain distribution-free prediction intervals. It can be considered as a combination of split conformal prediction (see Appendix A) and cross-validation. It works as follows: data are divided into $K$ disjoint subsets (or folds) $I_{1},\dots,I_{K}$ of size $m=n/K$. The (cross-validation) scores are defined as:
@@
This implies that the statement in (4) can be proved by combining the results in (6) and (7). For a detailed discussion on cross-conformal prediction see, for example, Chapter 4.4 of Vovk et al., 2022a .
@@ 3.2 Modified cross-conformal prediction
Since p-values take discrete values, in order to avoid having noninformative sets identical to $\mathcal{Y}$, the inequality $1<\alpha(m+1)$ must hold. A slightly improvement can be obtained using randomized p-values $P_{1}(Y_{n+1};\tau),\dots,P_{K}(Y_{n+1};\tau)$ defined by
$\begin{split}P_{k}(y;\tau)=\frac{\tau+\sum_{i\in I_{k}}\tau\mathds{1}\left\{s\left((X_{n+1},y);\mathcal{D}_{[n]\setminus I_{k}}\right)=S_{i}^{\mathrm{CV}}\right\}+\sum_{i\in I_{k}}\mathds{1}\left\{s\left((X_{n+1},y);\mathcal{D}_{[n]\setminus I_{k}}\right)<S_{i}^{\mathrm{CV}}\right\}}{m+1}\end{split}$ (9)
@@ 4 New variants of cross-conformal prediction
In this section, we improve the prediction sets in (8) using recent results regarding the combination of p-values. In particular, the combination rules that will be used are more powerful than the combinations valid under arbitrary dependence of the p-values. The results are obtained in a completely general manner and do not require the use of expensive computational procedures (Carlsson et al.,, 2014) or the use of specific models (Boström et al.,, 2017).
@@ 4.1 Exchangeable modified cross-conformal prediction
Let $P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1})$ be the (cross-conformal) p-values obtained using data $Z_{i}=(X_{i},Y_{i}),\,i=[n+1],$ then $P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1})$ are exchangeable, meaning that $\mathbf{P}\stackrel{{\scriptstyle d}}{{=}}\mathbf{P}^{\pi}$, where =$\stackrel{{\scriptstyle d}}{{=}}$ represents equality in distribution, $\mathbf{P}=(P_{1}(Y_{n+1}),\dots,P_{K}(Y_{n+1}))$, $\mathbf{P}^{\pi}=(P_{\pi(1)}(Y_{n+1}),\dots,$ $P_{\pi(K)}(Y_{n+1}))$ and $\pi:[K]\to[K]$ is any permutation of the indices.
@@
It holds that $\hat{C}_{n,K,\alpha}^{\mathrm{e\text{-}mod\text{-}cross}}(X_{n+1})\subseteq\hat{C}_{n,K,\alpha}^{\mathrm{mod\text{-}cross}}(X_{n+1})$. In addition, if data are exchangeable,
@@
Once $K$ exchangeable (or more generally dependent) p-values are obtained, there are several methods to combine them. The proposed solution is to use the minimum (over $\ell$) of the mean obtained using the first $\ell$ p-values, which is related to the valid combination rule “twice the average” used by Vovk et al., (2018); Vovk et al., 2022a to prove the coverage guarantee of cross-conformal prediction. However, similar results apply to other merging functions like quantiles (for example “twice the median” is also a valid combination rule) and generalized averages (e.g., geometric mean or harmonic mean). However, it is important to emphasize that, as explained in Section 6 of Vovk and Wang, (2020), the mean is an effective method of combining strongly dependent p-values, which is often the case for rank-based p-values obtained by cross-conformal prediction. Other merging rules, such as the Bonferroni method, are more appropriate near independence. For instance, Lei et al., (2018, Section 2.3) shows, under some assumptions, that the Bonferroni rule is overly conservative in the context of multisplit conformal prediction.
@@ 4.2 Randomized modified cross-conformal prediction
In the previous paragraph, we leveraged the exchangeability of p-values to obtain a smaller set. In this section, we move in a different direction and improve the set (8) using a simple “randomization trick” (introducing a uniform random variable). Indeed, in this case, the exchangeability of the p-values is not necessary. As before, the improvement does not alter the marginal validity of the set, but the new result is obtained in a different way. Although randomization is avoided in some statistical applications due to the extra randomness it introduces, in this case, it does not pose a major issue. Indeed, cross-conformal prediction is, by definition, a randomized method. More precisely, data are randomly divided into $K$ different subsets in the first step, which means that the procedure inherently includes randomness (see Remark 4.8 for further discussion).
@@
$\begin{split}\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})=\Bigg{\{}y\in\mathcal{Y}:\frac{1}{2-U}\frac{1}{K}\sum_{k=1}^{K}P_{k}(y)>\alpha\Bigg{\}},\end{split}$ (12)
@@
It holds that $\hat{C}^{\mathrm{u\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$. In addition, if data are exchangeable,
@@ 4.3 Exchangeable and randomized modified cross-conformal prediction
$\begin{split}\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})=\Bigg{\{}y\in\mathcal{Y}:\min\left\{\frac{1}{2-U}P_{1}(y),\min_{\ell\in[K]}\frac{1}{\ell}\sum_{k=1}^{\ell}P_{k}(y)\right\}>\alpha\Bigg{\}},\end{split}$ (14)
@@
It holds that $\hat{C}^{\mathrm{eu\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{e\text{-}mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})\subseteq\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha}(X_{n+1})$. In addition, if data are exchangeable,
@@
The set in (14) can be considered an improvement of the set described in (10) but not of the (randomized) set in (12), since only the first p-value of the sequence is randomized.
@@
A direct use of external randomization is present in both procedures described in Section 4.2 and Section 4.3. The use of randomization is often avoided in statistical methods, as it can pose challenges to the reproducibility of results. Clearly, randomization becomes problematic when a human is in the loop and runs the procedure multiple times until the desired result is achieved (for example, in the described cases, one can sample $U$ many times until it reaches a value close to zero). Some recommendations aimed at solving this problem are proposed, for example, in Ramdas and Manole, (2025, Section 10). Actually, in the data pipeline of split and cross-conformal prediction methods, randomization comes into play in different parts: by default in the division of data into their respective folds; to smoothen p-values as described in (9); and potentially to improve the conditional coverage as described in Hore and Barber, (2024). In particular, there exists a trade-off between reproducibility and statistical efficiency, and it is not always evident which should be prioritized. In other words, randomized procedures tend to be more efficient than standard procedures but may lack in terms of reproducibility, and vice versa. For instance, our methods may be particularly well-suited in industrial settings, where hundreds or thousands of predictions are made daily, and efficiency may be more important.